At December 31, 2022, Swifty Corporation reported the following plant assets.

Land

$ 4,353,000

Buildings

$26,720,000

Less: Accumulated depreciation—buildings

17,303,175

9,416,825

Equipment

58,040,000

Less: Accumulated depreciation—equipment

7,255,000

50,785,000

Total plant assets

$64,554,825


During 2023, the following selected cash transactions occurred.

Apr. 1 Purchased land for $3,192,200.
May 1 Sold equipment that cost $870,600 when purchased on January 1, 2016. The equipment was sold for $246,670.
June 1 Sold land for $2,321,600. The land cost $1,451,000.
July 1 Purchased equipment for $1,596,100.
Dec. 31 Retired equipment that cost $1,015,700 when purchased on December 31, 2013. No salvage value was received.

Journalize the transactions. Swifty uses straight-line depreciation for buildings and equipment. The buildings are estimated to have a 40-year useful life and no salvage value; the equipment is estimated to have a 10-year useful life and no salvage value. Update depreciation on assets disposed of at the time of sale or retirement. (Record entries in the order displayed in the problem statement. Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

Record adjusting entries for depreciation for 2023. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)


Prepare the plant assets section of Swifty’s balance sheet at December 31, 2023. (Hint: You may wish to set up T accounts, post beginning balances, and then post 2023 transactions.) (List Plant Assets in order of Land, Building and Equipment.)
Answer #1
a) Date Account titles and explanation Debit Credit
Apr 1. Land 3192200
Cash 3192200
(Purchased Land)
May 1. Depreciation expense (Note:1) 29020
Accumulated depreciation-Equipment 29020
(Depreciation recorded prior to sale)
May 1. Cash 246670
Accumulated depreciation-equipment (Note:1) 638440
Equipment 870600
Gain on sale of equipment (Balancing figure) 14510
(Disposal of equipment)
June 1. Cash 2321600
Land 1451000
Gain on sale of Land (Balancing figure) 870600
(Sale of Land)
July 1. Equipment 1596100
Cash 1596100
(Purchased equipment)
Dec 31. Depreciation expense (Note:2) 101570
Accumulated depreciation-Equipment 101570
(Depreciation recorded prior to retirement)
Dec 31. Accumulated depreciation-equipment 1015700
Equipment 1015700
(Retired equipment and no proceeds were received)
Note:1- Accumulated depreciation of equipment sold:
Original cost=$ 870600
Acquired on Jan 1,2016
Useful life=10 years
Depreciation per year=(Original cost-Salvage value)/Useful life=(870600-0)/10=$ 87060
Depreciation will be same from 2009 till disposal since straight line method is used
Depreciation from 2016 to 2022=87060*7=$ 609420
Depreciation for 2023:
From Jan 1 to May 1 (For 4 months)=87060*(4/12)=$ 29020
Accumulated depreciation=609420+29020=$ 638440
Note:2- Accumulated depreciation of equipment retired:
Original cost=$ 1015700
Acquired on Dec 31,2013
Useful life=10 years
Depreciation per year=(Original cost-Salvage value)/Useful life=(1015700-0)/10=$ 101570
Depreciation will be same from 2014 till disposal since straight line method is used
Depreciation from 2014 to 2022=101570*9=$ 914130
Depreciation for 2023=$ 101570
Accumulated depreciation=914130+101570=$ 1015700
b) Adjusting entry:
Date Account titles and explanation Debit Credit
Dec 31. Depreciation expense 668000
Accumulated depreciation-Building 668000
(Note:3)
(Depreciation recorded for building)
Depreciation expense 5774980
Accumulated depreciation-Equipment 5774980
(Note:4)
(Depreciation recorded for equipment)
Note:3
Original cost=$ 26720000
Useful life=40 years
Depreciation per year=(Original cost-Salvage value)/Useful life=(26720000-0)/40=$ 668000
Note:4
Depreciation of equipment:
Cost of equipment eligible for 12 months depreciation and depreciation expense:
$
Beginning balance 58040000
Less:
Original cost of equipment sold 870600
Original cost of equipment retired 1015700 1886300
Cost of equipment eligible for 12 months depreciation 56153700
Useful life in years 10
Depreciation expense 5615370
Cost of equipment eligible for 6 months depreciation and depreciation expense:
Cost of equipment purchased on July 1=$ 1596100
Useful life=10 years
Depreciation expense for 6 months=1596100/10=$ 159610
Total cost of equipment=56153700+1596100=$ 57749800
Total depreciation=5615370+159610=$ 5774980
c) Property,Plant and equipment
$ $
Land (4353000+3192200-1451000) 6094200
Building 26720000
Less: Accumulated depreciation (17303175+668000) 17971175 8748825
Equipment (Refer note:4) 57749800
Less: Accumulated depreciation (7255000+5774980) 13029980 44719820
Total plant assets 59562845